MATH 375 – Probability and Statistics 1 Second Group Assignments

Group 1:

• Ben Daly

• Robert (Packy) Loughran

• Katherine McGovern

• Kathryn Noonan

Group 2:

• Becca Bates

• Ingrid Gliottone

• Luke Jaroszewicz

• Tom Potvin

Group 3:

• Erin Lord

• Kaitlin Mearn

• Lauren Pereira

• Courtney Walls

Group 4:

• Rosie Arcuri

• Michelle Hedrick

• Kaitlin Horndahl

• Jackie Varanelli

Group 5:

• Gerry Bueti

• Jackie Markot

• Beth Richard

• Sheena Stangler

Group 6:

• Erin Cagianello

• Maggie Kielman

• Angelina Kim

• Galo Lopez

Group 7:

• Natalie Cate

• Tasia Peña

• John Restivo

• Christina Roberto